# Global parameterizations of $\pi\pi$ scattering with dispersive constraints: beyond the S0 wave

### v1

@@ IV.1.5 Improved matching with the high-energy regime
Following [70, 73, 71, 72, 54], the input for the FDRs at high energies is a relatively simple parameterization of data on $NN$, $\pi N$ and $\pi\pi$ total cross sections (see [95] for the formalism, and [54] for the updated treatment). It is important to remark that Regge theory is just a semi-local approach expected to provide an average description, particularly once resonances start to appear. Therefore, at any given energy, it is not expected to match exactly the amplitude reconstructed from partial waves. Since the interest of [70, 73, 71, 54, 72, 96] was mostly in the low-energy region or in resonances up to 1 GeV (external $s$ variable in Eqs. (IV.1.1) and (57)), the Regge regime was set to start at 1.42 GeV for all FDRs (internal $s^{\prime}$ variable in Eqs. (IV.1.1) and (57)). As seen in Fig. 17, no particular attention was paid to provide a smooth matching between the partial-wave reconstruction of $\textrm{Im}F^{i}$ below 1.42 GeV and the Regge regime above. Actually, the matching in the $F^{0+}$ case was pretty bad (green versus blue at 1.42 GeV). Such discontinuities in the integrand could give rise to nearby artifacts in the FDR output, as seen in Fig. 15. Of course, with the focus in the region below 1.1 GeV, this was not a big concern for [54]. However, in this work, we aim to obtain a precise description up to energies of 1.41.41.4\,1.4GeV and above.
Figure 17: $\mathrm{Im}\,F^{00},\mathrm{Im}\,F^{0+}$ and $\mathrm{Im}\,F^{I_{t}=1}$ in the 1.2 to 1.8 GeV region. Since Ref. [54] was focused on energies below 1.1 GeV, a crude matching between the CFD and the Regge description at 1.42 GeV (blue-dotted vertical line) was considered good enough. For a precise description up to higher energies, we now use as input the Global Fits up to 1.62 GeV (black-dotted vertical line) for $F^{0+}$ and $F^{I_{t}=1}$, and Regge above. This matching point displacement clearly improves the matching with the Regge parameterization. Note that, for safety, we only consider the FDRs up to 20 MeV below the matching point (continuous versus dotted vertical lines). We use Global Fit I for illustration, but the situation is similar for Global Fits II and III.
For this reason, we have decided to move higher the matching point with the Regge regime; from 1.42 GeV to the round energy value close to the point where the $F^{i}$ amplitudes, reconstructed as a sum of partial waves, match the Regge description within uncertainties. These plots justify why the input is matched with the Regge description at 1.42, 1.62, and 1.62 GeV for the $F^{00},F^{0+}$ and $F^{I_{t}=1}$ FDRs, respectively. Still, as also done [54], for safety we only use the output of FDRs up to an energy about 20 MeV below the matching point. Namely, the FDRs are applied as constraints only up to 1.4, 1.6, and 1.6 GeV, respectively.
@@
In summary, by choosing a better matching point between the amplitude reconstructed from partial waves and the Regge description, we hope to have soothed any concern about the appearance of unphysical artifacts. Moreover, it has allowed us to impose two FDRs up to 200 MeV above their previous application.
@@ IV.2 Roy-like dispersion relations
Roy equations for $\pi\pi$ scattering [75] (see [50, 51, 52, 53, 54, 45, 49] for recent analyses and reviews) are partial-wave dispersion relations where crossing symmetry has been used to rewrite the left-cut (crossed-channel) contributions in terms of s-channel partial waves. Thus, they form a coupled system for the infinite tower of partial waves. At sufficiently low energies, one can concentrate on those with the lowest angular momentum, typically, the S0, S2, and P waves (and occasionally, D and F waves, see [55]). In such a case, higher partial waves and high energy information are provided as input and gathered in the so-called driving terms. In their most common form, which we use here, their derivation implies the integration in $t$ of fixed-$t$ dispersion relations to obtain partial-wave dispersion relations. This means their applicability range is limited to an energy of around $1.14$ GeV. The original Roy equations were derived with two subtractions, but there is also a once-subtracted version called GKPY equations [54]. We refer to both of them as “Roy-like” equations. As we will see below, this different number of subtractions implies that, when calculated from the same data without further theoretical input, Roy equations are more precise than GKPY equations at low energies and vice versa at high energies. This is because Roy uncertainties grow rapidly with energy, whereas GKPY uncertainties do not grow so fast. In practice, Roy uncertainties are smaller than those of GKPY below the resonance region, i.e., $\lesssim 0.5{\mathrm{\,Ge\kern-1.00006ptV}}$, and larger above. Hence, Roy equations are particularly powerful in constraining threshold parameters [50, 51] or the low-energy constants of ChPT. Actually, when supplemented with ChPT constraints, Roy equations can be solved for the S and P waves and provide predictions below some matching point, or even for the lightest resonances [50, 51, 52, 97]. However, here we simply adhere to a data-driven formalism and use the dispersion relations as constraints on the fits. Nevertheless, we refer to  [50, 51, 52, 53, 54, 45, 49] for explicit expressions and further details.
@@
In the previous section, we imposed Forward Dispersion Relations to obtain constrained Global Fits from the threshold to 1.4 GeV. Recall that below 0.9 GeV we are mimicking the CFD parameterizations of [54]. They were obtained as data fits constrained to satisfy the three FDRs, as well as Roy and GKPY equations for the S0, P, and S2 waves. The only significant improvements in the present work happen above 0.9 GeV, and, for several waves, these changes are small until 1.4 GeV. Consequently, Roy and GKPY equations are already very well satisfied within uncertainties by just imposing FDRs. The only exception is the S2 GKPY equation, whose output comes about 1.5 deviations away above 1 GeV for the Global Fits constrained only with FDRs. This mismatch requires a small fix in our Global Fits.
To that end, we have also imposed Roy and GKPY equations together with FDRs, starting from the Global Fits obtained after minimizing the FDRs first. The changes are very small but amend the small deviation in the S2 GKPY equation. As a matter of fact, all the plots for the constrained Global Fit that we have provided before are fully constrained with three FDRs, three Roy, and three GKPY equations. In Fig. 18 we illustrate the nice fulfillment of Roy and GKPY equations by the Global Fit I. In those plots, we can see that, as discussed above, the Roy equation uncertainties are smaller than those of GKPY at low energies but larger in the resonance region. We can also see that both Roy and GKPY equations are well satisfied within errors. The only point slightly beyond 1 deviation is the highest one at 1.114 GeV, for the GKPY S2 wave.
@@
Table 9: Average distances of Roy and GKPY equations for the three Global Fits. These “Roy-like” dispersion relations, reaching only up to 1.1 GeV, are remarkably well described within uncertainties for the three fits.
The situation is very similar for Global Fits II and III. Actually, we provide in Table 9 the values of the average distances (defined as in Eq. (58)) between output and input of Roy and GKPY equations. The three Global Fits satisfy remarkably well all Roy-like dispersion relations in all regions of interest.
@@ V Summary and Discussion
In this work, we have provided Global Fits, which are relatively simple sets of dispersively constrained $\pi\pi\rightarrow\pi\pi$ fits to data on scattering partial waves. Specifically, we describe the P, S2, D0, D2, F, G0, and G2 waves. The S0 wave parameterization is not revisited here because it was already studied in great detail in [68] together with the P wave that we have improved here above $\sim 0.9\,$GeV. We only discuss in an appendix the small changes in the S0 parameters induced by the larger changes in the other waves.
@@
By “Global” we mean that the parameterizations are continuous in the whole energy region where they are defined, and also have a continuous derivative, except as required by thresholds. Moreover, they extend from the $\pi\pi$ threshold up to 1.8 GeV or somewhat beyond, depending on where the data on each partial wave cease to exist. They are dispersively constrained because, when fitting the data we impose as penalty functions, the averaged distances between the direct and dispersive calculation of nine dispersion relations. These are: three Forward Dispersion Relations (FDRs) for the full amplitude as well as three Roy equations and three GKPY equations (one each for the S0, P, and S2 partial waves). Roy and GKPY equations are used up to 1.1 GeV, which is their applicability limit. The $\pi^{0}\pi^{0}$ FDR, dominated by the S0 wave, is applied up to 1.4 GeV as was done in [68]. In addition, we have extended up to 1.6 GeV the two FDRs that involve the P wave.
@@
To this end, we have revisited the P-wave parameterization, slightly updating the input in the elastic region with the recent pion form factor data analysis in [77], and improving the description above 0.9 GeV. The latter has been achieved by allowing its inelasticity to open at the $\pi\omega$ threshold and using better parameterizations that do not make the uncertainty grow artificially with the energy. This wave has suffered the largest update, but the parameterizations of other waves have also been modified to become inelastic somewhat below the $K\bar{K}$ threshold and to have more uniform, smaller, and more realistic uncertainties. Moreover, we have also provided new parameterizations for the F and G0 waves, to reproduce their resonant behavior, as well as the G2 wave, since they are needed to test FDRs up to higher energies than before. For this same reason, we have extended the description of all the waves beyond 1.4 GeV, often with more flexible parameterizations. As a result, we have improved the accuracy of our partial waves and reduced the uncertainty band in the dispersion relations output. Consequently, a better matching with the Regge description than in previous works is required. This is achieved by increasing the matching point, which also allows us to extend the FDR applicability region. Simultaneously, the stability of the fits is improved and the uncertainties near the matching point become smaller. All in all, the new constrained Global Fits satisfy very well the FDRs even though their uncertainty bands are considerably smaller, particularly for the two relations involving the P wave.
@@
Finally, let us note that we have studied separately three different sets of data, called Solutions I, II, and III here, which give rise to three Global Fits. Their S2, D2, and G waves as well as all other waves up to energies of around $0.9{\mathrm{\,Ge\kern-1.00006ptV}}$ are almost indistinguishable, which means that the three of them satisfy Roy and GKPY equations very well. Above that energy, the situation changes from wave to wave. For the P wave, the three fits are qualitatively similar but incompatible up to 1.4 GeV, where they start to differ widely. For the F wave, they are qualitatively similar but incompatible within their uncertainties in the whole energy region. The three fits for the D0 wave are indistinguishable for the phase shift up to 1.3 GeV but the elasticity is very different from one another from 0.9 GeV.
@@
The original data in Solutions II and III had very tiny or lacked uncertainties in some waves. Even assuming errors similar to Solution I, the average FDR fulfillment is somewhat worse and much less homogeneous for Global Fits II and III. In particular, these two fits have two regions $\sim$100 MeV wide, where it has not been possible to make them satisfy the $I_{t}=1$ FDR within uncertainties. The first of those regions, sitting around 1 GeV, favors the onset of the D0-wave inelasticity at $\sim$0.9 GeV instead of the original onset at the $K\bar{K}$ threshold. In addition, the S0-wave inelasticity of Solutions II and III prefers the “non-dip” scenario above 1 GeV, known to be disfavored by GKPY equations, although we have made them follow the “dip” scenario. Moreover, when FDRs are imposed, Global Fits II and III deviate somewhat further from their original data set than Global Fit I. None of these single caveats is enough, by itself, to discard either Global Fit II or III, and that is why we have provided them together with Global Fit I. Nevertheless, Global Fit I, free of these caveats, seems to be slightly favored against the other two, and we have used it to illustrate our constrained fitting procedure.
@@ Acknowledgements.
We thank, G. Colangelo, C. Hanhart, M. Hoferichter, B. Kubis, and P. Stoffer for discussions and comments, particularly pointing out possible improvements upon previous works by JRP, JRE, and collaborators. We are particularly indebted to G. Colangelo, M. Hoferichter, and P. Stoffer for providing their results for the pion vector form factor. JRP thanks the kind hospitality of the Helmholtz-Institut für Strahlen-und Kernphysik (Theorie), Universität Bonn, Germany, where he spent three months, while working on this manuscript, as an invited visiting researcher under the MCIN “Salvador de Madariaga” Program Grant PRX22/00129. This work is part of the Grant PID2022-136510NB-C31 funded by MCIN/AEI/ 10.13039/501100011033, it has also received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No.824093. P. R. is supported by the MIU (Ministerio de Universidades, Spain) fellowship FPU21/03878, and J.R.E. by the Ramón y Cajal program (RYC2019-027605-I) of the Spanish MINECO.

### v2

@@ IV.1.5 Improved matching with the high-energy regime
Figure 17: $\mathrm{Im}\,F^{00},\mathrm{Im}\,F^{0+}$ and $\mathrm{Im}\,F^{I_{t}=1}$ in the 1.2 to 1.8 GeV region. Since Ref. [54] was focused on energies below 1.1 GeV, a crude matching between the CFD and the Regge description at 1.42 GeV (blue-dotted vertical line) was considered good enough. For a precise description up to higher energies, we now use as input the Global Fits up to 1.62 GeV (black-dotted vertical line) for $F^{0+}$ and $F^{I_{t}=1}$, and Regge above. This matching point displacement clearly improves the matching with the Regge parametrization. Note that, for safety, we only consider as constraints the FDRs up to 20 MeV below the matching point (continuous versus dotted vertical lines). We use Global Fit I for illustration, but the situation is similar for Global Fits II and III.
Following [70, 73, 71, 72, 54], the input for the FDRs at high energies is a relatively simple parametrization of data on $NN$, $\pi N$ and $\pi\pi$ total cross sections (see [99] for the formalism, and [54] for the updated treatment). It is important to remark that Regge theory is just a semi-local approach expected to provide an average description, particularly once resonances start to appear. Therefore, at any given energy, it is not expected to match exactly the amplitude reconstructed from partial waves. Since the interest of [70, 73, 71, 54, 72, 100] was mostly in the low-energy region or in resonances up to 1 GeV (external $s$ variable in Eqs. (IV.1.1) and (57)), the Regge regime was set to start at 1.42 GeV for all FDRs (internal $s^{\prime}$ variable in Eqs. (IV.1.1) and (57)). As seen in Fig. 17, no particular attention was paid to provide a smooth matching between the partial-wave reconstruction of $\textrm{Im}F^{i}$ below 1.42 GeV and the Regge regime above. Actually, the matching in the $F^{0+}$ case was pretty bad (green versus blue at 1.42 GeV). Such discontinuities in the integrand could give rise to nearby artifacts in the FDR output, as seen in Fig. 15 for the CFD. Of course, with the focus in the region below 1.1 GeV, this was not a big concern for [54]. However, in this work, we aim to obtain a precise description up to energies of 1.41.41.4\,1.4GeV and above.
For this reason, we have decided to move higher the matching point with the Regge regime; from 1.42 GeV to the round energy value close to the point where the $F^{i}$ amplitudes, reconstructed as a sum of partial waves, match the Regge description within uncertainties. These plots justify why the partial-wave input is matched with the Regge description at 1.42, 1.62, and 1.62 GeV for the $F^{00},F^{0+}$ and $F^{I_{t}=1}$ FDRs, respectively. Still, as also done [54], for safety we only use the output of FDRs up to an energy about 20 MeV below the matching point. Namely, the FDRs are applied as constraints only up to 1.4, 1.6, and 1.6 GeV, respectively.
@@
In summary, by choosing a better matching point between the amplitude reconstructed from partial waves and the Regge description, we hope to have soothed any concern about the appearance of unphysical artifacts. Moreover, it has allowed us to impose two FDRs up to 200 MeV above their previous region of application.
@@ IV.2 Roy-like dispersion relations
Roy equations for $\pi\pi$ scattering [75] (see [50, 51, 52, 53, 54, 45, 49] for recent analyses and reviews) are partial-wave dispersion relations where crossing symmetry has been used to rewrite the left-cut (crossed-channel) contributions in terms of s-channel partial waves. Thus, they form a coupled system for the infinite tower of partial waves. At sufficiently low energies, one can concentrate on those with the lowest angular momentum, typically, the S0, S2, and P waves (and occasionally, D and F waves, see [55]). In such a case, higher partial waves and high energy information are provided as input and gathered in the so-called driving terms. In their most common form, which we use here, their derivation implies the integration in $t$ of fixed-$t$ dispersion relations to obtain partial-wave dispersion relations. This means their applicability range is limited to an energy of around $1.14$ GeV. The original Roy equations were derived with two subtractions, but there is also a once-subtracted version called GKPY equations [54]. We refer to both of them as “Roy-like” equations. As we will see below, this different number of subtractions implies that, when calculated from the same data without further theoretical input, Roy equations are more precise than GKPY equations at low energies and vice versa at high energies. This is because Roy uncertainties grow rapidly with energy, whereas GKPY uncertainties do not grow so fast. In practice, Roy uncertainties are smaller than those of GKPY below the resonance region, i.e., $s^{1/2}\lesssim 0.5{\mathrm{\,Ge\kern-1.00006ptV}}$, and larger above. Hence, Roy equations are particularly powerful in constraining threshold parameters [50, 51] or the low-energy constants of ChPT. Actually, when supplemented with ChPT constraints, Roy equations can be solved for the S and P waves and provide predictions below some matching point, or even for the lightest resonances [50, 51, 52, 101]. However, here we simply adhere to a data-driven formalism and use the dispersion relations as constraints on the fits. Nevertheless, we refer to  [50, 51, 52, 53, 54, 45, 49] for explicit expressions and further details.
@@
In the previous section, we imposed Forward Dispersion Relations to obtain constrained Global Fits from the threshold to 1.4 GeV. Recall that below 0.9 GeV we are mimicking the CFD parametrizations of [54]. They were obtained as data fits constrained to satisfy the three FDRs, as well as Roy and GKPY equations for the S0, P, and S2 waves. The only significant improvements in the present work happen above 0.9 GeV, and, for several waves, these changes are small until 1.4 GeV. Consequently, Roy and GKPY equations are already very well satisfied within uncertainties by just imposing FDRs. The only exception is the S2 GKPY equation above 1 GeV, whose output comes about 1.5 deviations away for the Global Fits constrained only with FDRs. This mismatch requires a small fix in our Global Fits.
@@
Table 9: Average quadratic distances of Roy and GKPY equations for the three Global Fits. These “Roy-like” dispersion relations, reaching only up to 1.1 GeV, are remarkably well described within uncertainties for the three fits.
To that end, we have also imposed Roy and GKPY equations together with FDRs, starting from the Global Fits obtained after minimizing the FDRs first. The changes are very small but amend the small deviation in the S2 GKPY equation. As a matter of fact, all the plots for the constrained Global Fits that we have provided before are fully constrained with three FDRs, three Roy, and three GKPY equations. In Fig. 18 we illustrate the nice fulfillment of Roy and GKPY equations by the Global Fit I. In those plots, we can see that, as discussed above, the Roy equation uncertainties are smaller than those of GKPY at low energies but larger in the resonance region. We can also see that both Roy and GKPY equations are well satisfied within errors. The only point slightly beyond 1 deviation is the highest one at 1.114 GeV, for the GKPY S2 wave.
The situation is very similar for Global Fits II and III. Actually, we provide in Table 9 the values of the average quadratic distances (defined as in Eq. (58)) between output and input of Roy and GKPY equations. The three Global Fits satisfy remarkably well all Roy-like dispersion relations in all regions of interest.
@@ V Summary and Discussion
In this work, we have provided Global Fits, which are relatively simple sets of dispersively constrained $\pi\pi\rightarrow\pi\pi$ fits to data on scattering partial waves. Specifically, we describe the P, S2, D0, D2, F, G0, and G2 waves. The S0 wave parametrization is not revisited here because it was already studied in great detail in [68] together with the P wave that we have improved here above $\sim 0.9\,$GeV. We only discuss in an appendix the small changes in the S0 parameters induced by the larger changes in the other waves.
@@
By “Global” we mean that the parametrizations are continuous in the whole energy region where they are defined, and also have a continuous derivative, except as required by thresholds. Moreover, they extend from the $\pi\pi$ threshold up to 1.8 GeV or somewhat beyond, depending on where the data on each partial wave cease to exist. They are dispersively constrained because, when fitting the data, we impose as penalty functions the averaged quadratic distances between the direct and dispersive calculation of nine dispersion relations. These are: three Forward Dispersion Relations (FDRs) for the full amplitude as well as three Roy equations and three GKPY equations (one each for the S0, P, and S2 partial waves). Roy and GKPY equations are used up to 1.1 GeV, which is their applicability limit. The $\pi^{0}\pi^{0}$ FDR, dominated by the S0 wave, is applied up to 1.4 GeV as was done in [68]. In addition, we have extended up to 1.6 GeV the two FDRs that involve the P wave.
@@
To this end, we have revisited the P-wave parametrization, slightly updating the input in the elastic region with the recent pion form factor data analysis in [77], and improving the description above 0.9 GeV. The latter has been achieved by allowing its inelasticity to open at the $\pi\omega$ threshold and using better parametrizations that do not make the uncertainty grow artificially with the energy. This wave has suffered the largest update, but the parametrizations of other waves have also been modified to become inelastic somewhat below the $K\bar{K}$ threshold and to have more uniform, smaller, and more realistic uncertainties. Moreover, we have also provided new parametrizations for the F and G0 waves, to reproduce their resonant behavior, as well as the G2 wave, since they are needed to test FDRs up to higher energies than before. For this same reason, we have extended the description of all the waves beyond 1.4 GeV, often with more flexible parametrizations. As a result, we have improved the accuracy of our partial waves and reduced the uncertainty band in the dispersion relations output. Consequently, a better matching with the Regge description than in previous works is required. This is achieved by increasing the matching point, which also allows us to extend the FDR applicability region. Simultaneously, the stability of the fits is improved and the uncertainties near the matching point become smaller. All in all, the new constrained Global Fits satisfy very well the FDRs even though their uncertainty bands are considerably smaller, particularly for the two relations involving the P wave.
@@
In addition, let us note that we have studied separately three different sets of data, called Solutions I, II, and III here, which give rise to three Global Fits. Their S2, D2, and G waves as well as all other waves up to energies of around $0.9{\mathrm{\,Ge\kern-1.00006ptV}}$ are almost indistinguishable, which means that the three of them satisfy Roy and GKPY equations very well. Above that energy, the situation changes from wave to wave. For the P wave, the three fits are qualitatively similar but incompatible up to 1.4 GeV, where they start to differ widely. For the F wave, they are qualitatively similar but incompatible within their uncertainties in the whole energy region. The three fits for the D0 wave are indistinguishable for the phase shift up to 1.3 GeV but the elasticity is very different from one another from 0.9 GeV.
@@
The original data in Solutions II and III had very tiny or lacked uncertainties in some waves. Even assuming errors similar to Solution I, the average FDR fulfillment is somewhat worse and much less homogeneous for Global Fits II and III. In particular, these two fits have two regions $\sim$100 MeV wide, where it has not been possible to make them satisfy the $I_{t}=1$ FDR within uncertainties. The first of those regions, sitting around 1 GeV, favors the onset of the D0-wave inelasticity at $\sim$0.9 GeV instead of the original onset at the $K\bar{K}$ threshold. In addition, the S0-wave elasticity of Solutions II and III prefers the “non-dip” scenario above 1 GeV, known to be disfavored by GKPY equations, although we have made them follow the “dip” scenario. Moreover, when FDRs are imposed, Global Fits II and III deviate somewhat further from their original data set than Global Fit I. None of these single caveats is enough, by itself, to discard either Global Fit II or III. That is why we have provided them together with Global Fit I. Nevertheless, Global Fit I, free of these caveats, seems to be slightly favored against the other two, and we have used it to illustrate our constrained fitting procedure.
Finally, it is important to emphasize that all the nominal resonance parameters in our parametrizations, ( like $\hat{m}_{\rho}$, $m_{f_{2}}$, $m_{\rho_{3}}$, $m_{f_{4}}$, $\Gamma_{\rho_{3}}$, $\Gamma_{f_{4}}$, etc.), are merely parameters of phenomenological parametrizations intended for use on the real axis. Resonances are rigorously characterized by the position and residue of their associated $T$-matrix pole in the contiguous Riemann sheet. The naive extrapolation of our phenomenological formulas into the complex $s$-plane to determine resonance pole parameters is highly model-dependent. This model dependence can be eliminated—or at least significantly reduced—by employing proper analytic continuation techniques, such as partial-wave dispersion relations or other analytic continuation methods applied to the dispersive results on the real axis. Such analyses, however, lie beyond the scope of this manuscript and may be pursued in future work along the lines of Refs. [100, 69].
@@ Acknowledgements.
We thank, G. Colangelo, C. Hanhart, M. Hoferichter, B. Kubis, and P. Stoffer for discussions and comments, particularly pointing out possible improvements upon previous works by JRP, JRE, and collaborators. We are particularly indebted to G. Colangelo, M. Hoferichter, and P. Stoffer for providing their results for the pion vector form factor. JRP thanks the kind hospitality of the Helmholtz-Institut für Strahlen-und Kernphysik (Theorie), Universität Bonn, Germany, where he spent three months, while working on this manuscript, as an invited visiting researcher under the MCIN “Salvador de Madariaga” Program Grant PRX22/00129. This work is part of the Grant PID2022-136510NB-C31 funded by MCIN/AEI/ 10.13039/501100011033, it has also received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No.824093. P. R. is supported by the MIU (Ministerio de Universidades, Spain) fellowship FPU21/03878, and J.R.E. by the Ramón y Cajal program (RYC2019-027605-I) of the Spanish MICIU/AEI.
Appendix A Threshold parameters
Our aim in this work has been to obtain simple and global parametrizations of partial-wave data covering the energy range from threshold to at least 1.8 GeV, simultaneously consistent with the dispersive constraints. We have made them continuous and with a continuous derivative (except as required by thresholds).
Previous dispersive analyses were more focused on the low-energy region, particularly below 1.1 GeV, and, up to 0.9 GeV, we have basically mimicked them. Of course, to cover a much wider energy region with just one parametrization, and being just a mimic of the CFD at low energies, we have sacrificed the precision attained by the CFD. For the $\pi\pi$ threshold parameters, the old CFD is still more accurate, and even more when combined with sum rules, as done in [54].
Nevertheless, we have checked that our threshold parameters are consistent with the values in [54]. Thus, we have gathered in Table 10 the values of threshold parameters for the S, P, D, F, and G waves that result from our constrained Global Fits. We also provide the values used as input, obtained from the CFD parametrization in [54], except for those of the G waves, which are taken from [73]. In addition, we are listing the best values in [54]. These are the most reliable because they are not only obtained from the CFD but from the use of sum rules, and therefore have a much better precision and are parametrization independent.